# OfCaM: Global Human Mesh Recovery via Optimization-free Camera Motion Scale Calibration

### v1

@@ 1 Introduction
Human pose and shape estimation (also called Human Mesh Recovery, HMR) in world coordinates is a key component of many vision applications [24, 6]. There are many successful (local) HMR methods [10, 12, 15, 4, 17, 3, 20], but they work primarily in camera coordinates. Only a few world-coordinate HMR methods, i.e., global HMR, have been developed [38, 16, 29, 28]. Most of these approaches learn a local-to-global mapping directly from a sequence of 3D (local) meshes, as the mesh sequence themselves provides a strong cue. Yet in some cases, there is ambiguity when the background is ignored. Consider, for example, a person riding a skateboard vs. standing on the ground, both have a similar local motion  [^1] This work uses “motion” to refer to a sequence of human poses or meshes, or camera extrinsics over time. but totally different global motions (see Fig. 1(a)).
An observed (local) human mesh sequence is composed of the human motion in the global space relative to the camera motion. As such, global human motion can be formulated in terms of the local motion and the camera motion. Given that the camera-coordinate HMR is a mature area of research [17, 12, 4, 3, 20], decoupling the camera motion from the global motion is logical and feasible alternate solution [37, 14]. A typical approach to estimate camera motion is SLAM  [5, 31, 30]. SLAM relies on the identification and continuous tracking of static environmental reference points to establish a spatial map and compute the camera’s trajectory relative to these landmarks. One limitation of SLAM is that it only estimates camera motion up to an unknown scale factor. This is typically resolved in robotics applications by integrating additional sensors (e.g., Inertial Measurement Units) or using calibration tools (e.g., checkerboard) to establish a metric scale. However, these solutions are not directly applicable to arbitrary videos for human motion analysis.
Therefore, some recent global HMR methods [37, 14] attempt to solve for the SLAM scale factor through optimization. The optimization, based on a loss function that evaluates the consistency between 3D human meshes projection and 2D video evidence, alongside smoothness constraints, jointly solves for the scale, human mesh, and camera trajectory. However, the inherent entanglement of human and camera motion makes it very challenging, sometimes leading to scale estimations that are off by a factor of several times, exemplified by the discrepancies between SLAHMR [37]’s and the ground true trajectory in Fig. 1(b). Another drawback is that the optimizations is very time-consuming; processing a one-minute video takes several minutes or longer (See Fig. 1(c)).
@@
Figure 1: (a) Video sequence as an entanglement of the camera and human motion in the world coordinate. (b) and (c) Regression-based methods like WHAM [28] are time-efficient but fail in ambiguous cases; Optimization-based methods like SLAHMR [37] struggle to optimize a good trajectory and are time-consuming; while ours can achieve accurate trajectory and optimization-free.
SLAM works well when reference points from the static background can be tracked. A typical setting in which SLAM fails is when the moving foreground takes up a majority of the scene [1, 2], which may happen when humans are too close to the camera. To that end, we design a SLAM failure indicator and revert to a local-to-global human motion mapping to compensate when it indicates the failure as the local-to-global mapping depends less on the background information.
@@ 2.1 World Coordinate Human Mesh Recovery
Image-based HMR methods [10, 33, 36, 22, 26, 19, 18, 25, 27, 3] traditionally focus on recovering human meshes within the camera’s coordinate system. Although major video-based HMR advancements [12, 4, 20] also operate within same camera space, the advent of video data has paved the way for HMR exploration in world coordinates [38, 16, 9, 37, 29, 28]. The transition to world coordinates introduces the distinct challenge of disentangling both dynamic camera motion and human motion, which is a relatively nascent research area. While most previous attempts (e.g., GLAMR [38], DnD [16], TRACE [29], and WHAM [28]) proposed to infer global motion from observable local behaviors (e.g., if a person looks like they are walking, it is assumed they are moving forwards globally), the inherent ambiguities of local-to-global mapping present significant limitations. In contrast, our approach does not solely depend on these dataset-learned local-to-global priors but rather employs them as additional cues to enhance accuracy.
Recent efforts including SLAHMR [37] and PACE [14] recognize the utility of background information in determining camera motion with SLAM techniques [23, 30, 31]. However, these methods aim to address the ‘unknown scale’ problem in SLAM outputs by jointly optimizing scale, pose, and shape parameters—a procedure that is inherently ambiguous. Our method, by contrast, deviates from these intensive optimization strategies by calibrating the scale factor directly, utilizing the depth predictions from HMR models, and thereby giving an optimization-free solution.
@@ 2.2 Camera Calibration
Camera calibration is a fundamental procedure in robotics and computer vision that enables precise spatial measurement and scene reconstruction. Typically, this process relies on additional sensors, such as Inertial Measurement Units (IMUs) [39, 8], or reference markers like checkerboards [7] to define a known metric scale. However, these traditional calibration methods are not feasible for arbitrary human-centric videos within the HMR domain, due to the absence of external sensors and standardized calibration tools. In contrast, single-view metrology [40] suggests that objects with well-defined geometrical priors can themselves act as natural calibration references. Motivated by this concept, HMR models, with their inherent human geometric priors, have the potential to serve as surrogate calibration devices. In our work, we utilize the predicted absolute joint depths from the HMR model as the reference to accurately and efficiently calibrate the unknown scale factor.
@@ 3.1 Human Motion in Camera Coordinates $\mathcal{M}_{c}$
The 3D human motion from a video $I={\{I_{t}\}}_{t=1}^{T}$ of $T$ frames can be represented in the camera space by a $T$-length sequence of SMPL parameters $\mathcal{M}_{c}={\{\boldsymbol{\theta}_{t},\boldsymbol{\beta}_{t},\boldsymbol{\psi}_{t},\boldsymbol{\tau}_{t}\}}_{t=1}^{T}$. SMPL [21] is a widely used 3D statistical model of the human body. For a given frame at time $t$, the SMPL model maps body pose $\boldsymbol{\theta}_{t}\in\mathbb{R}^{23\times 3}$, shape parameters $\boldsymbol{\beta}_{t}\in\mathbb{R}^{10}$, root orientation $\boldsymbol{\psi}_{t}\in\mathbb{R}^{3}$, and root translation $\boldsymbol{\tau}_{t}\in\mathbb{R}^{3}$ to a 3D mesh of the human body $\mathbf{V}_{t}\!\in\!\mathbb{R}^{6890\times 3}$ in the camera space. The individual joints can be mapped from the SMPL parameters with the function $\mathbf{J}=\mathcal{J}(\boldsymbol{\theta}_{t},\boldsymbol{\beta}_{t},\boldsymbol{\psi}_{t},\boldsymbol{\tau}_{t})\in\mathbb{R}^{24\times 3}$.
Note that $\boldsymbol{\psi}_{t}$ and $\boldsymbol{\tau}_{t}$ are sometimes referred to as “global” orientation and translation parameters by the SMPL model. However, HMR models [10, 15, 12, 4, 13, 35] estimate these parameters with respect to the camera extrinsics $\mathcal{E}$ at frame $t$, where =$\mathcal{E}={\{{\boldsymbol{R}}_{t},{\boldsymbol{T}}_{t}\}}_{t=1}^{T}$ is the sequence of camera rotations ${\boldsymbol{R}}_{t}\in\mathbb{R}^{3\times 3}$ and translations ${\boldsymbol{T}}_{t}\in\mathbb{R}^{3}$.
@@ 3.2 Human Motion in World Coordinates $\mathcal{M}_{w}$
This equation explains how we decouple the global motion from the local estimation by isolating and removing the camera motion, which serves as one of our key insights. Nevertheless, getting the correct camera extrinsics $\mathcal{E}$ can be difficult. SLAM is a widely used method to estimate camera motion [23, 30, 31], though it can only predict the camera extrinsics ${\{\boldsymbol{R}_{t},s\cdot\boldsymbol{T}_{t}\}}_{t=1}^{T}$ up to an unknown scale $s$. Our work focus on how to calibrate the scale $s$ based on the recovered human mesh.
@@ 4 Method
Our pipeline is illustrated in Fig. 2. For global HMR, the inputs are typically captured by moving cameras featuring static background content and dynamic foreground content of humans. Prior approaches [38, 16, 29] infer global human motion $\mathcal{M}_{w}$ exclusively from foreground’s local motion $\mathcal{M}_{c}$. In contrast, recent attempts [37, 14] jointly optimize global human motion $\mathcal{M}_{w}$ and the SLAM dereived camera motion $\mathcal{E}$ to fit with the 2D observation. Different from their complex optimization, we propose an optimization-free way to calibrate the scale by comparing the depth of some key reference points from the output of SLAM and HMR, where we select the human-background contact joints as the reference points. (Sec. 4.2). Furthermore, to resolve the cases of problematic SLAM output, we introduce the global human motion refinement via fusing local motion priors (Sec. 4.2).
@@ 4.2 Local-to-global Motion Adjustment
Local-to-global Human Motion. The previous works [38, 16, 29] in global HMR are devoted to learning the global motion from local motion priors. In our work, we adopt a lightweight sub-module from GLAMR [38] as the local-to-global motion predictor. The input is a sequence of the body pose $\{\boldsymbol{\theta}_{t}\}$ and the output is the sequence of human’s global orientation $\{\tilde{\boldsymbol{\Phi}}_{t}\}$ and global translation $\{\tilde{\boldsymbol{\Gamma}}_{t}\}$ both relative to the first frame. Thus, we can get the global motion derived by local-to-global predictor:
@@
Segment-wise Adaptive Global Motion Fusion. Since when SLAM failed, the scale is no longer reliable but may retain potentially useful shape information, we opt to fuse the local-to-global motion $\mathcal{M}^{L}_{w}(\hat{I})$ with SLAM-derived motion $\mathcal{M}^{S}_{w}(\hat{I})$ as the final motion for failed segments $\hat{I}\in\mathcal{S}$. Thus, we first align the SLAM-derived motion to the local-to-global motion by Umeyama’s method [32]: $U(\mathcal{M}^{S}_{w}(\hat{I}),\mathcal{M}^{L}_{w}(\hat{I}))={\{\boldsymbol{\theta}_{t},\ \boldsymbol{\beta}_{t},\ {\boldsymbol{\Phi}}^{\prime}_{t},\ {\boldsymbol{\Gamma}}^{\prime}_{t}\}}_{t\in\hat{I}}$ where $U(a,b)$ is Umeyama’s method to align points set $a$ to points set $b$. The fused global motion is the weighted average of these two motions:
@@ 5.1 Implementation Details, Dataset and Metrics
Our experiments  [^3] The code will be released upon acceptance adopt DPVO [31] as the SLAM model for the camera motion stream and CLIFF [17] as the HMR model for the human motion stream. The distance threshold $\delta=400px$ for outlier rejection and the standard deviation threshold $\upsilon=2$ for SLAM failure segment identification.
Datasets. Following previous works [28], we evaluate the global human motion and camera motion on a subset of EMDB  [11] (EMDB 2), which contains 25 sequences captured by the dynamic camera and provides ground truth global motion for both human and camera.
Metrics for Human Motion. Same with previous works [14, 37, 28], we evaluate human’s global motion error by: (1) WA-MPJPE which is the average Euclidean distance between the ground truth and the predicted joint positions (i.e., MPJPE) after aligning each segment for every 100 frames; (2) W-MPJPE which is the MPJPE error after only aligning the first two frames of each 100-frames segments with the ground truth. (3) RTE which is the human’s root translation error of the whole sequence after the rigid alignment. We also evaluate local mesh error by (4) PA-MPJPE which is the MPJPE error after Procrustes aligned with ground truth.
Metrics for Camera Motion. We follow SLAM convention and previous works [14] for camera motion evaluation, reporting (1) ATE which is the Average Translation Error after rigidly aligning the camera trajectories; (2) ATE-S which is Average Translation Error without Scale alignment, providing a more accurate reflection of inaccuracies in the captured scale of the scene.
@@ 5.2 Ablation Study and Analysis
GLAMR [38] 56.0 726.6 280.8 16.7
TRACE [29] 58.0 1702.3 529.0 18.9
SLAHMR [37] 61.5 776.1 326.9 10.2
WHAM(w/ DPVO) [28] 41.9 354.8 135.6 6.0

### v2

@@ 1 Introduction
Human pose and shape estimation (also called Human Mesh Recovery, HMR) in world coordinates is a key component of many vision applications [AR_rauschnabel2022what, VR_han2022virtual]. There are many successful (local) HMR methods [HMR_2017, VIBE_2020, SPIN_2019, TCMR_2021, CLIFF_2022, ImpHMR_2023, FAMI_2022], but they work primarily in camera coordinates. Only a few world-coordinate HMR methods, i.e., global HMR, have been developed [GLAMR_2022, DnD_2022, TRACE_2023, WHAM_2023]. Most of these approaches learn a local-to-global mapping directly from a sequence of 3D (local) meshes, as the mesh sequence themselves provides a strong cue. Yet in some cases, there is ambiguity when the background is ignored. Consider, for example, a person riding a skateboard vs. standing on the ground, both have a similar local motion  [^1] This work uses “motion” to refer to a sequence of human poses or meshes, or camera extrinsics over time. but totally different global motions (see Fig. 1(a)).
An observed (local) human mesh sequence is composed of the human motion in the global space relative to the camera motion. As such, global human motion can be formulated in terms of the local motion and the camera motion. Given that the camera-coordinate HMR is a mature area of research [CLIFF_2022, VIBE_2020, TCMR_2021, ImpHMR_2023, FAMI_2022], decoupling the camera motion from the global motion is logical and feasible alternate solution [SLAMHR_2023, PACE_2023]. A typical approach to estimate camera motion is SLAM  [SLAM_2006, DPVO_2023, DROIDSLAM_2021]. SLAM relies on the identification and continuous tracking of static environmental reference points to establish a spatial map and compute the camera’s trajectory relative to these landmarks. One limitation of SLAM is that it only estimates camera motion up to an unknown scale factor. This is typically resolved in robotics applications by integrating additional sensors (e.g., Inertial Measurement Units) or using calibration tools (e.g., checkerboard) to establish a metric scale. However, these solutions are not directly applicable to arbitrary videos for human motion analysis.
Therefore, some recent global HMR methods [SLAMHR_2023, PACE_2023] attempt to solve for the SLAM scale factor through optimization. The optimization, based on a loss function that evaluates the consistency between 3D human meshes projection and 2D video evidence, alongside smoothness constraints, jointly solves for the scale, human mesh, and camera trajectory. However, the inherent entanglement of human and camera motion makes it very challenging, sometimes leading to scale estimations that are off by a factor of several times, exemplified by the discrepancies between SLAHMR [SLAMHR_2023]’s and the ground true trajectory in Fig. 1(b). Another drawback is that the optimizations is very time-consuming; processing a one-minute video takes several minutes or longer (See Fig. 1(c)).
@@
Figure 1: (a) Video sequence as an entanglement of the camera and human motion in the world coordinate. (b) and (c) Regression-based methods like WHAM [WHAM_2023] are time-efficient but fail in ambiguous cases; Optimization-based methods like SLAHMR [SLAMHR_2023] struggle to optimize a good trajectory and are time-consuming; while ours can achieve accurate trajectory and optimization-free.
SLAM works well when reference points from the static background can be tracked. A typical setting in which SLAM fails is when the moving foreground takes up a majority of the scene [DynamicSLAM_2018, SLAMsurvey_2016], which may happen when humans are too close to the camera. To that end, we design a SLAM failure indicator and revert to a local-to-global human motion mapping to compensate when it indicates the failure as the local-to-global mapping depends less on the background information.
@@ 2.1 World Coordinate Human Mesh Recovery
Image-based HMR methods [HMR_2017, BodyNet_2018, DenseBody_2019, I2l-meshnet_2020, pifuhd_2020, lintransformers_2021, Meshgraphormer_2021, rong2019delving, sengupta2020synthetic, ImpHMR_2023] traditionally focus on recovering human meshes within the camera’s coordinate system. Although major video-based HMR advancements [VIBE_2020, TCMR_2021, FAMI_2022] also operate within same camera space, the advent of video data has paved the way for HMR exploration in world coordinates [GLAMR_2022, DnD_2022, bodyslam_2022, SLAMHR_2023, TRACE_2023, WHAM_2023]. The transition to world coordinates introduces the distinct challenge of disentangling both dynamic camera motion and human motion, which is a relatively nascent research area. While most previous attempts (e.g., GLAMR [GLAMR_2022], DnD [DnD_2022], TRACE [TRACE_2023], and WHAM [WHAM_2023]) proposed to infer global motion from observable local behaviors (e.g., if a person looks like they are walking, it is assumed they are moving forwards globally), the inherent ambiguities of local-to-global mapping present significant limitations. In contrast, our approach does not solely depend on these dataset-learned local-to-global priors but rather employs them as additional cues to enhance accuracy.
Recent efforts including SLAHMR [SLAMHR_2023] and PACE [PACE_2023] recognize the utility of background information in determining camera motion with SLAM techniques [ORBSLAM_2015, DROIDSLAM_2021, DPVO_2023]. However, these methods aim to address the ‘unknown scale’ problem in SLAM outputs by jointly optimizing scale, pose, and shape parameters—a procedure that is inherently ambiguous. Our method, by contrast, deviates from these intensive optimization strategies by calibrating the scale factor directly, utilizing the depth predictions from HMR models, and thereby giving an optimization-free solution.
@@ 2.2 Camera Calibration
Camera calibration is a fundamental procedure in robotics and computer vision that enables precise spatial measurement and scene reconstruction. Typically, this process relies on additional sensors, such as Inertial Measurement Units (IMUs) [zhang2000flexible, heikkila1997four], or reference markers like checkerboards [hartley2003multiple] to define a known metric scale. However, these traditional calibration methods are not feasible for arbitrary human-centric videos within the HMR domain, due to the absence of external sensors and standardized calibration tools. In contrast, single-view metrology [singleviewmetrology_2020] suggests that objects with well-defined geometrical priors can themselves act as natural calibration references. Motivated by this concept, HMR models, with their inherent human geometric priors, have the potential to serve as surrogate calibration devices. In our work, we utilize the predicted absolute joint depths from the HMR model as the reference to accurately and efficiently calibrate the unknown scale factor.
@@ 3.1 Human Motion in Camera Coordinates $\mathcal{M}_{c}$
The 3D human motion from a video $I={\{I_{t}\}}_{t=1}^{T}$ of $T$ frames can be represented in the camera space by a $T$-length sequence of SMPL parameters $\mathcal{M}_{c}={\{\boldsymbol{\theta}_{t},\boldsymbol{\beta}_{t},\boldsymbol{\psi}_{t},\boldsymbol{\tau}_{t}\}}_{t=1}^{T}$. SMPL [SMPL_2015] is a widely used 3D statistical model of the human body. For a given frame at time $t$, the SMPL model maps body pose $\boldsymbol{\theta}_{t}\in\mathbb{R}^{23\times 3}$, shape parameters $\boldsymbol{\beta}_{t}\in\mathbb{R}^{10}$, root orientation $\boldsymbol{\psi}_{t}\in\mathbb{R}^{3}$, and root translation $\boldsymbol{\tau}_{t}\in\mathbb{R}^{3}$ to a 3D mesh of the human body $\mathbf{V}_{t}\!\in\!\mathbb{R}^{6890\times 3}$ in the camera space. The individual joints can be mapped from the SMPL parameters with the function $\mathbf{J}=\mathcal{J}(\boldsymbol{\theta}_{t},\boldsymbol{\beta}_{t},\boldsymbol{\psi}_{t},\boldsymbol{\tau}_{t})\in\mathbb{R}^{24\times 3}$.
Note that $\boldsymbol{\psi}_{t}$ and $\boldsymbol{\tau}_{t}$ are sometimes referred to as “global” orientation and translation parameters by the SMPL model. However, HMR models [HMR_2017, SPIN_2019, VIBE_2020, TCMR_2021, PARE_2021, vitpose_2022] estimate these parameters with respect to the camera extrinsics $\mathcal{E}$ at frame $t$, where =$\mathcal{E}={\{{\boldsymbol{R}}_{t},{\boldsymbol{T}}_{t}\}}_{t=1}^{T}$ is the sequence of camera rotations ${\boldsymbol{R}}_{t}\in\mathbb{R}^{3\times 3}$ and translations ${\boldsymbol{T}}_{t}\in\mathbb{R}^{3}$.
@@ 3.2 Human Motion in World Coordinates $\mathcal{M}_{w}$
This equation explains how we decouple the global motion from the local estimation by isolating and removing the camera motion, which serves as one of our key insights. Nevertheless, getting the correct camera extrinsics $\mathcal{E}$ can be difficult. SLAM is a widely used method to estimate camera motion [ORBSLAM_2015, DROIDSLAM_2021, DPVO_2023], though it can only predict the camera extrinsics ${\{\boldsymbol{R}_{t},s\cdot\boldsymbol{T}_{t}\}}_{t=1}^{T}$ up to an unknown scale $s$. Our work focus on how to calibrate the scale $s$ based on the recovered human mesh.
@@ 4 Method
Our pipeline is illustrated in Fig. 2. For global HMR, the inputs are typically captured by moving cameras featuring static background content and dynamic foreground content of humans. Prior approaches [GLAMR_2022, DnD_2022, TRACE_2023] infer global human motion $\mathcal{M}_{w}$ exclusively from foreground’s local motion $\mathcal{M}_{c}$. In contrast, recent attempts [SLAMHR_2023, PACE_2023] jointly optimize global human motion $\mathcal{M}_{w}$ and the SLAM dereived camera motion $\mathcal{E}$ to fit with the 2D observation. Different from their complex optimization, we propose an optimization-free way to calibrate the scale by comparing the depth of some key reference points from the output of SLAM and HMR, where we select the human-background contact joints as the reference points. (Sec. 4.2). Furthermore, to resolve the cases of problematic SLAM output, we introduce the global human motion refinement via fusing local motion priors (Sec. 4.2).
@@ 4.2 Local-to-global Motion Adjustment
Local-to-global Human Motion. The previous works [GLAMR_2022, DnD_2022, TRACE_2023] in global HMR are devoted to learning the global motion from local motion priors. In our work, we adopt a lightweight sub-module from GLAMR [GLAMR_2022] as the local-to-global motion predictor. The input is a sequence of the body pose $\{\boldsymbol{\theta}_{t}\}$ and the output is the sequence of human’s global orientation $\{\tilde{\boldsymbol{\Phi}}_{t}\}$ and global translation $\{\tilde{\boldsymbol{\Gamma}}_{t}\}$ both relative to the first frame. Thus, we can get the global motion derived by local-to-global predictor:
@@
Segment-wise Adaptive Global Motion Fusion. Since when SLAM failed, the scale is no longer reliable but may retain potentially useful shape information, we opt to fuse the local-to-global motion $\mathcal{M}^{L}_{w}(\hat{I})$ with SLAM-derived motion $\mathcal{M}^{S}_{w}(\hat{I})$ as the final motion for failed segments $\hat{I}\in\mathcal{S}$. Thus, we first align the SLAM-derived motion to the local-to-global motion by Umeyama’s method [umeyama_1991]: $U(\mathcal{M}^{S}_{w}(\hat{I}),\mathcal{M}^{L}_{w}(\hat{I}))={\{\boldsymbol{\theta}_{t},\ \boldsymbol{\beta}_{t},\ {\boldsymbol{\Phi}}^{\prime}_{t},\ {\boldsymbol{\Gamma}}^{\prime}_{t}\}}_{t\in\hat{I}}$ where $U(a,b)$ is Umeyama’s method to align points set $a$ to points set $b$. The fused global motion is the weighted average of these two motions:
@@ 5.1 Implementation Details, Dataset and Metrics
Our experiments  [^3] The code will be released upon acceptance adopt DPVO [DPVO_2023] as the SLAM model for the camera motion stream and CLIFF [CLIFF_2022] as the HMR model for the human motion stream. The distance threshold $\delta=400px$ for outlier rejection and the standard deviation threshold $\upsilon=2$ for SLAM failure segment identification.
Datasets. Following previous works [WHAM_2023], we evaluate the global human motion and camera motion on a subset of EMDB  [EMDB_2023] (EMDB 2), which contains 25 sequences captured by the dynamic camera and provides ground truth global motion for both human and camera.
Metrics for Human Motion. Same with previous works [PACE_2023, SLAMHR_2023, WHAM_2023], we evaluate human’s global motion error by: (1) WA-MPJPE which is the average Euclidean distance between the ground truth and the predicted joint positions (i.e., MPJPE) after aligning each segment for every 100 frames; (2) W-MPJPE which is the MPJPE error after only aligning the first two frames of each 100-frames segments with the ground truth. (3) RTE which is the human’s root translation error of the whole sequence after the rigid alignment. We also evaluate local mesh error by (4) PA-MPJPE which is the MPJPE error after Procrustes aligned with ground truth.
Metrics for Camera Motion. We follow SLAM convention and previous works [PACE_2023] for camera motion evaluation, reporting (1) ATE which is the Average Translation Error after rigidly aligning the camera trajectories; (2) ATE-S which is Average Translation Error without Scale alignment, providing a more accurate reflection of inaccuracies in the captured scale of the scene.
@@ 5.2 Ablation Study and Analysis
GLAMR [GLAMR_2022] 56.0 726.6 280.8 16.7
TRACE [TRACE_2023] 58.0 1702.3 529.0 18.9
SLAHMR [SLAMHR_2023] 61.5 776.1 326.9 10.2
WHAM(w/ DPVO) [WHAM_2023] 41.9 354.8 135.6 6.0